# AI-Assisted Rapid Crystal Structure Generation Towards a Target Local Environment

### v1

@@ Abstract
In the field of material design, traditional crystal structure prediction approaches require extensive structural sampling through computationally expensive energy minimization methods using either force fields or quantum mechanical simulations. While emerging artificial intelligence (AI) generative models have shown great promise in generating realistic crystal structures more rapidly, most existing models fail to account for the unique symmetries and periodicity of crystalline materials, and they are limited to handling structures with only a few tens of atoms per unit cell. Here, we present a symmetry-informed AI generative approach called Local Environment Geometry-Oriented Crystal Generator (LEGO-xtal) that overcomes these limitations. Our method generates initial structures using AI models trained on an augmented small dataset, and then optimizes them using machine learning structure descriptors rather than traditional energy-based optimization. We demonstrate the effectiveness of LEGO-xtal by expanding from 25 known low-energy sp2 carbon allotropes to over 1,700, all within 0.5 eV/atom of the ground-state energy of graphite. This framework offers a generalizable strategy for the targeted design of materials with modular building blocks, such as metal–organic frameworks and next-generation battery materials.
@@ I INTRODUCTION
Recently, the advent of artificial intelligence (AI) generative models has opened up new avenues for rapid crystal structure generation. Inspired by the successes of synthetic image generation and AI-driven protein structure prediction Jumper et al. (2021), various AI-based generative models have been introduced to the materials community since 2018, including Generative Adversarial Networks (GANs) Kim et al. (2020); Zhu et al. (2024), Variational Autoencoders (VAEs) Noh et al. (2019), and more recent active learning Merchant et al. (2023), diffusion Xie et al. (2021); Jiao et al. (2023); Zeni et al. (2025); Levy et al. (2024); Jiao et al. (2024); Miller et al. (2024); Zhong et al. (2025) and transformer models Antunes, Butler, and Grau-Crespo (2024); Gruver et al. (2024); Cao et al. (2024); Kazeev et al. (2025). These models offer a fundamentally new approach to exploring structure space by learning data-driven representations, allowing for the efficient generation of low-energy crystal structures at a much faster rate than traditional global optimization methods. However, the direct application of AI-based models to the CSP task faces a significant challenge: the generated structures often lack diversity or fail to satisfy essential physical constraints, as these models tend to either replicate structural prototypes from the training data, or generate invalid and unstable structure, rather than producing truly novel and physical-plausible structures Cheetham and Seshadri (2024). This limitation arises because these data-driven models primarily generate crystal structures only based on the statistical patterns from the training data source without incorporating the unique characteristics of crystalline materials, particularly their symmetries. To address the challenge, several recent works have attempted to incorporate symmetry into generative AI model design with the aim of enhancing the quality and validity of generated structures Jiao et al. (2024); Levy et al. (2024); Antunes, Butler, and Grau-Crespo (2024); Cao et al. (2024); Kazeev et al. (2025); Zhu et al. (2024). However, most of the previous works simply add each individual symmetry representation into the existing AI models. Consequently, the learned models may be able to generate new candidate structures with a symmetry identical to that in the training samples. Nevertheless, they are less likely to generate those closely symmetry-related structures due to the distinct symmetry representations, thus largely limiting the model’s learning efficiency.
@@
In this work, we present a novel symmetry-informed generative model called LEGO-xtal (Local Environment Geometry-Oriented Crystal Generator) that overcomes the limitations of existing AI models in crystal structure prediction. Our approach is designed to generate realistic crystal structures with well-defined local environments (i.e. specified chemical building blocks) while preserving the inherent symmetries and periodicity characteristic of crystalline materials. As a proof of concept, we train our generative model on an augmented dataset of known three-coordinated carbon allotropes —commonly referred to as sp2 carbon structures— which enables the exploration of a broad range of structural configurations. Starting from only 25 known low-energy sp2 carbon allotropes (as well as additional 115 high-energy examples), we demonstrate the effectiveness of our approach by generating over 1,700 new structures with varied dimensionality, topology and unit cell sizes.
@@ II.1 Symmetry-based Crystal Representation
Thus, one essentially needs two sets of variables to describe a structure: (i) the discrete variables denoting the exact space group symmetry and Wyckoff site indices; and (ii) the continuous variables describing the cell parameters and Wyckoff positions. For high symmetry cases, the cell parameters and $(x,y,z)$ coordinates may have additional constraints. For instance, cubic unit cells must have $a=b=c$ and $\alpha=\beta=\gamma=90^{\circ}$. The Wyckoff position may be completely or partially restricted to specific fractional coordinates due to site symmetry requirements. If input variables do not satisfy symmetry requirements, deterministic algorithms can check if variables can be symmetrized within an acceptable threshold. Considering the symmetry constraints, the free continuous variables are thus reduced to a 1-length vector [3.567].
Therefore, we can use a fixed-length 1D vector to represent a crystal structure with a maximum number of Wyckoff positions. For a maximum of 8 Wyckoff positions, the vector length is 1 + 6 + 8$\times$5 = 46. For a single-component system, this reduces to 39.
@@ II.2 Practical Challenges of AI Generative Models for Crystal Structure Prediction
Using this representation, the objective of crystal structure prediction is to find or generate the best combination of these variables that leads to a low-energy configuration. Inspired by the successes of AI generative models for image, point cloud, or text, we aim to develop a generative model to learn the statistical distribution from the training data for both discrete and continuous variables, and then generate new tabular data to represent novel crystal structures. However, there are several major challenges that prohibit the direct application of existing AI generative models to crystal structure generation.
First, there is a risk of lacking sufficient training data to describe the design space for the discrete variables (e.g., space group and Wyckoff site choices). In the context of crystal structure prediction, most materials generation tasks require the researchers to generate likely candidates from only a limited choices of good examples. However, with a total of 230 space groups and 1731 Wyckoff site choices, even for a simple unit cell of 10 atoms, the design space becomes astronomically large. With limited training data (typically ranging from several dozen to ten thousands examples), models are more likely to struggle to learn the underlying distribution of crystal structures. This lack of diversity and coverage can lead to overfitting and poor generalization, resulting in generated structures that are too similar to the training data or fail to explore the design space effectively.
Second, assuming the discrete design space (i.e. space group and Wyckoff site choices) has been sufficiently covered, the next challenge is to generate a valid crystal structure under the symmetry constraints. Currently, there exist two common practices to generate the continuous variables: (i) directly applying the diffusion model to generate the continuous variables for unit cell and Wyckoff sites Levy et al. (2024); Jiao et al. (2024); or (ii) combining random sampling with energy-based optimization to search for the low-energy structures Kazeev et al. (2025); Zhu et al. (2024). The first approach still relies on the availability of sufficient training data and may not be applied to small data tasks. The second approach is more reliable. However, for a specific space group and Wyckoff site choice, we may need to sample a large number of random configurations and then optimize them with the objective of finding the nearest lowest-energy configuration. This is often time-consuming and the quality of the optimized solution is overly reliant on the initial guess.
@@ II.3 Symmetry and Geometry Considerations
Pre-relaxation with respect to the reference local environment. Once the model is capable of suggesting a promising candidate crystal, we then need to optimize the remaining continuous variables to get the realistic structure, under the conditioned space group symmetry constraint. In many practical applications, we may already know how the atoms may behave in a short-range cutoff. For instance, carbon may take sp, sp2, or sp3 local bonding; Porous materials such as zeolites and metal–organic frameworks (MOFs) are often categorized by their well-defined molecular building units, making them ideal targets for generative approaches that preserve local chemical environments. Therefore, we can utilize this knowledge to predefine the reference environment according to the recently developed power spectrum descriptors from spherical harmonics expansion Steinhardt, Nelson, and Ronchetti (1983); Kazhdan, Funkhouser, and Rusinkiewicz (2003); Bartók, Kondor, and Csányi (2013); Zagaceta, Yanxon, and Zhu (2020); Yanxon et al. (2020), and then relax the structure with respect to this reference environment. Figure 2b explains the idea of using SO(3) descriptor based reference environment for a randomly generated symmetric crystal. The numerical details can be found in the supplementary materials. Compared to the commonly used energy-based optimization, it not only saves a lot of computational time, but also effectively further reduces the searching space by excluding the possibility of handling low-energy configurations with an undesired local environment. Hence, we can rapidly screen if the trial structure can be relaxed to the desired environment prior to further consideration with more expensive energy-based optimization. Note that similar ideas of optimizing the structures at the descriptor space have also been explored in several recent works Tao, Shao, and Zhu (2024); Pickard (2025).
@@ III.1 Training Data Augmentation & Tabularization
First, we need to gather the already known data used for training as many as possible. For carbon research, the online Samara Carbon Allotrope Database (SACADA) database, which contains all available data about 3-periodic carbon allotropes extracted from scientific literature from Web of Science and Scopus databases Hoffmann et al. (2016). From SACADA, we extracted 140 carbon allotropes which are made of all sp2 carbon atoms. This includes a total of 126 unique (space group, Wyckoff sites) combinations, with energy spanning from 0 (graphite) to 1.344 eV/atom using the GGA-PBE pseudopotential Perdew, Burke, and Ernzerhof (1996) within the framework of density functional theory (DFT).
For each structure, we utilized PyXtal’s subgroup symmetry module Fredericks et al. (2021) to generate valid subgroup representations within constraints of 500 atoms and 8 Wyckoff sites per unit cell. Since the subgroup representation often results in larger unit cells and more Wyckoff sites, we excluded 36 structures with either low symmetry or large atom counts from the original 140 structures. For each of the remaining structures, we generated 2-15 subgroup symmetry representations. Additionally, we randomly sampled the multiplicity of each Wyckoff site due to symmetry operations. To evaluate the effectiveness of data augmentation, we created three datasets labeled as v1-60k, v2-120k, and v3-240k, with details summarized in Table 1.
@@
Dataset   Number of samples Number of unique symmetries
@@ III.2 Generative Model Training & Sampling
Prior to employing AI generative models, we transformed each row into an extended feature representation to enhance model learning. For discrete features, one-hot encoding is used, where each unique categorical value is mapped to an individual binary column. In this representation, a value of 1 indicates the presence of the corresponding category, while all other columns are set to 0. For continuous columns, we apply a Bayesian Gaussian Mixture Model (BGMM) approach as proposed by Patki et al Patki, Wedge, and Veeramachaneni (2016), in which each continuous value is first scaled to the interval $[-1,1]$, and then softly assigned to distinct Gaussian components via a probabilistic one-hot representation. This dual encoding strategy effectively captures both normalized numerical information and probabilistic cluster associations, improving the stability and efficacy of feature representation during generative model training. In the dis-dataset, we replace the BGMM-based encoding with fixed binning. All continuous features are discretized into categorical bins (e.g., cell lengths into 100 bins between 0–35 Å), and the resulting values are one-hot encoded. This yields a fully categorical representation for the entire dataset.
After transformation, the table expands to 525 columns in the cont-dataset and 2339 columns in the dis-dataset. These two encoding schemes allow us to benchmark model performance under both mixed-type and fully discrete data conditions. Next, we explored two generative models on both dis- and cont-type datasets to learn the underlying distribution of crystal symmetry data. The models are trained to synthesize new rows of tabular features that correspond to novel crystal structures that satisfy crystallographic constraints.
@@
For more details regarding GAN and VAE models, please refer to the Supplementary Materials Section 1 and Figs. S1 and S2). In a typical experiment, we train each GAN or VAE model with 250-1000 epochs, which is sufficient to converge the loss function values. Then we take the epoch that achieves the lowest loss function value to generate 100k - 500k samples for further consideration.
@@ III.3 Structure Relaxation to Targeted Local Environment
To represent the local atomic environment, the SO(3) descriptor $P$, with the option to include the explicit radial distribution function, is calculated to extract the radial and angular features of the neighbor density function Yanxon et al. (2020). For the purpose of this work, we selected a sp2 carbon (i.e., one centered atom is connected to 3 neighboring carbon with 1.42 Å  bond length and 120 bond angles in the same plane) to compute its power spectrum $P_{\text{ref.}}$ up to a radial cutoff of 2.0 Å. For each generated structure, we first evaluated the corresponding power spectrum $P$ for each atom, and then sought to minimize the summed mean squared error (MSE) loss function as follows:
@@
During optimization, symmetry constraints were naturally enforced (due to the use of reduced variables) to ensure chemical realism and structural validity. Combining with the subgroup representation, this descriptor-guided strategy not only shortens the computational overhead associated with iterative energy evaluations but also allows the system to escape undesirable local minima associated with conventional force-based optimization. For more details, please refer to Fig. S3 and a motivating example in the Supplementary Section.

### v2

@@ Abstract
In the field of material design, traditional crystal structure prediction approaches require structural sampling through computationally expensive energy minimization methods using either force fields or quantum mechanical simulations. While emerging artificial intelligence (AI) generative models have shown great promise in generating realistic crystal structures more rapidly, most existing models fail to account for the unique symmetries and periodicity of crystalline materials, and they are limited to handling structures with only a few tens of atoms per unit cell. Here, we present a symmetry-informed AI generative approach called Local Environment Geometry-Oriented Crystal Generator (LEGO-xtal) that overcomes these limitations. Our method generates initial structures using AI models trained on an augmented small dataset, and then optimizes them using machine learning structure descriptors rather than traditional energy-based optimization. We demonstrate the effectiveness of LEGO-xtal by expanding from 25 known low-energy sp2 carbon allotropes to over 1,700, all within 0.5 eV/atom of the ground-state energy of graphite. This framework offers a generalizable strategy for the targeted design of materials with modular building blocks, such as metal–organic frameworks and next-generation battery materials.
@@ I INTRODUCTION
Recently, the advent of artificial intelligence (AI) generative models has opened up new avenues for rapid crystal structure generation. Inspired by the successes of synthetic image generation and protein structure prediction Jumper et al. (2021), various AI-based generative models have been introduced to the materials community since 2018, including Generative Adversarial Networks (GANs) Kim et al. (2020); Zhu et al. (2024), Variational Autoencoders (VAEs) Noh et al. (2019), and more recent active learning Merchant et al. (2023), diffusion Xie et al. (2021); Jiao et al. (2023); Zeni et al. (2025); Levy et al. (2024); Jiao et al. (2024); Miller et al. (2024); Zhong et al. (2025) and transformer models Antunes, Butler, and Grau-Crespo (2024); Gruver et al. (2024); Cao et al. (2024); Kazeev et al. (2025). These models offer a fundamentally new approach to exploring structure space by learning data-driven representations, allowing for the efficient generation of low-energy crystal structures at a much faster rate than traditional global optimization methods. However, the direct application of AI-based models to the CSP task faces a significant challenge: the generated structures often lack diversity or fail to satisfy essential physical constraints, as these models tend to either replicate structural prototypes from the training data, or generate invalid and unstable structure, rather than producing truly novel and physical-plausible structures Cheetham and Seshadri (2024). This limitation arises because these data-driven models primarily generate crystal structures only based on the statistical patterns from the training data source without incorporating the unique characteristics of crystalline materials, particularly their symmetries. To address the challenge, several recent works have attempted to incorporate symmetry into generative AI model design with the aim of enhancing the quality and validity of generated structures Jiao et al. (2024); Levy et al. (2024); Antunes, Butler, and Grau-Crespo (2024); Cao et al. (2024); Kazeev et al. (2025); Zhu et al. (2024). However, most of the previous works simply add each individual symmetry representation into the existing AI models. Consequently, the learned models may be able to generate new candidate structures with a symmetry identical to that in the training samples. Nevertheless, they are less likely to generate those closely symmetry-related structures, thus largely limiting the model’s learning efficiency.
@@
In this work, we present a novel symmetry-informed generative model called LEGO-xtal (Local Environment Geometry-Oriented Crystal Generator) that overcomes the limitations of existing AI models in crystal structure prediction. Unlike the previous AI generative models that focus on generation of relatively small crystal systems from a large dataset without the explicit constraints on the local chemical environment, our approach is designed to generate complex crystal structures with well-defined local environments (i.e. specified chemical building blocks) while preserving the inherent symmetries and periodicity characteristic of crystalline materials. As a proof of concept, we train our generative models (VAEs and GANs) on an augmented dataset of known three-coordinated carbon allotropes —commonly referred to as sp2 carbon structures— which enables the exploration of a broad range of structural configurations. Starting from only 25 known low-energy sp2 carbon allotropes (as well as additional 115 high-energy examples), we demonstrate the effectiveness of our approach by generating over 1,700 new structures with varied dimensionality, topology and unit cell sizes.
@@ II.1 Symmetry-based Crystal Representation
Thus, one essentially needs two sets of variables to describe a structure: (i) the discrete variables denoting the exact space group symmetry and Wyckoff site indices; and (ii) the continuous variables describing the cell parameters and Wyckoff positions. For high symmetry cases, the cell parameters and $(x,y,z)$ coordinates may have additional constraints. For instance, cubic unit cells must have $a=b=c$ and $\alpha=\beta=\gamma=90^{\circ}$. The Wyckoff position may be completely or partially restricted to specific fractional coordinates due to site symmetry requirements. If input variables do not satisfy symmetry requirements, deterministic algorithms can check if variables can be symmetrized within an acceptable threshold. Considering the symmetry constraints, the free continuous variables are thus reduced to a 1-length vector [3.567] for the diamond carbon.
[…]
@@ II.2 Practical Challenges of AI Generative Models for Crystal Structure Prediction
Using this representation, the objective of CSP is to find or generate the best combination of these variables that leads to a low-energy configuration. Inspired by the successes of AI generative models for image, point cloud, or text, we aim to develop a model to learn the statistical distribution from the training data for both discrete and continuous variables, and then generate new crystal representations. However, there are several challenges that prohibit the direct application of existing AI generative models to crystal generation.
First, there is a risk of lacking sufficient training data to describe the design space for the discrete variables (e.g., space group and Wyckoff site choices). In the context of crystal structure prediction, most materials generation tasks require the researchers to generate likely candidates from only a limited choices of good examples. However, with a total of 230 space groups and 1731 Wyckoff site choices, even for a simple unit cell of 10 atoms, the design space becomes astronomically large. With limited training data in lack of diversity and coverage, it is likely to lead to overfitting and poor generalization, resulting in generated structures that are too similar to the training data or fail to explore the design space effectively.
Second, assuming the discrete design space (i.e. space group and Wyckoff site choices) has been sufficiently covered, the next challenge is to generate a valid crystal structure under the symmetry constraints. Currently, there exist two common practices to generate the continuous variables: (i) directly applying the diffusion model to generate the continuous variables for unit cell and Wyckoff sites Levy et al. (2024); Jiao et al. (2024); or (ii) combining random sampling with energy-based optimization to search for the low-energy structures Kazeev et al. (2025); Zhu et al. (2024). The first approach relies on the availability of sufficient training data and may not be applied to small data tasks. Although the second approach is more reliable, for a specific space group and Wyckoff site choice, we need to sample a large number of random configurations and then optimize to find the low-energy states. This is often time-consuming and the quality of the optimized solution is overly reliant on the initial guess.
Third, materials researchers may not only seek target structures optimized for a single metric such as energy or a specific property. In practice, materials are often designed to incorporate established structural motifs or to maintain a preferred local environment (e.g., specific coordination geometry). To our knowledge, existing generative models do not explicitly enforce local geometry during structure generation. As a result, generated structures may deviate from the desired local motifs, limiting their practical relevance for targeted materials design.
@@ II.3 Symmetry and Geometry Considerations
Pre-relaxation with respect to the reference local environment. Once the model is capable of suggesting a promising candidate crystal, we then need to optimize the remaining continuous variables to get the realistic structure, under the conditioned space group symmetry constraint. In many practical applications, we may already know how the atoms may behave in a short-range cutoff. For instance, carbon may take sp, sp2, or sp3 local bonding; Porous materials such as zeolites and metal–organic frameworks (MOFs) are often categorized by their well-defined molecular building units, making them ideal targets for generative approaches that preserve local chemical environments. Therefore, we can utilize this knowledge to predefine the reference environment according to the recently developed power spectrum descriptors from spherical harmonics expansion Steinhardt, Nelson, and Ronchetti (1983); Kazhdan, Funkhouser, and Rusinkiewicz (2003); Bartók, Kondor, and Csányi (2013); Zagaceta, Yanxon, and Zhu (2020); Yanxon et al. (2020), and then relax the structure with respect to this reference environment. Figure 2b explains the idea of using SO(3) descriptor based reference environment for a randomly generated symmetric crystal. Ideally, we aim to achieve two outcomes from the use of pre-relaxation, including (1) maintain the same chance if the initial guess is close to an ideal arrangement; and (2) enhance the chance of generating good structures even from a bad guess by pre-relaxation. The numerical details and justification can be found in the supplementary materials Section III. Hence, we can rapidly screen if the trial structure can be relaxed to the desired environment prior to further consideration with more expensive energy-based optimization. Note that similar ideas of optimizing the structures at the descriptor space have also been explored in several recent works Tao, Shao, and Zhu (2024); Pickard (2025).
@@ III.1 Training Data Augmentation & Tabularization
First, we need to gather the already known data used for training as many as possible. For carbon research, the online Samara Carbon Allotrope Database (SACADA) database, which contains all available data about 3-periodic carbon allotropes extracted from scientific literature from Web of Science and Scopus databases Hoffmann et al. (2016). From SACADA, we extracted 140 carbon allotropes which are made of all sp2 carbon atoms. This includes a total of 127 unique (space group, Wyckoff sites) combinations, with energy spanning from 0 (graphite) to 1.344 eV/atom using the GGA-PBE pseudopotential Perdew, Burke, and Ernzerhof (1996) within the framework of density functional theory (DFT).
For each structure, we utilized PyXtal’s subgroup symmetry module Fredericks et al. (2021) to generate valid subgroup representations within constraints of 500 atoms and 8 Wyckoff sites per unit cell. Since the subgroup representation often results in larger unit cells and more Wyckoff sites, we excluded 10 structures with either low symmetry or large atom counts from the original 150 structures (see Supplementary Materials Section IV-A and Table S2). For each of the remaining structures, we generated 2-15 subgroup symmetry representations. Additionally, we randomly sampled the multiplicity of each Wyckoff site due to symmetry operations. To evaluate the effectiveness of data augmentation, we created three datasets labeled as v1-60k, v2-120k, and v3-240k, with details summarized in Table 1.
@@
Dataset   Number of samples Number of unique symmetries [^1] The unique symmetry is defined according to the combination of space group and sorted list of Wyckoff indices. For instance, the diamond carbon is denoted as (227, 8a).
@@ III.2 Generative Model Training & Sampling
Prior to employing AI generative models, we transformed each row into an extended feature representation to enhance model learning. For discrete features, one-hot encoding is used, where each unique categorical value is mapped to an individual binary column. In this representation, a value of 1 indicates the presence of the corresponding category, while all other columns are set to 0. For continuous columns, we apply a Bayesian Gaussian Mixture Model (BGMM) approach as proposed by Patki et al Patki, Wedge, and Veeramachaneni (2016), in which each continuous value is first scaled to the interval [1,1][-1,1], and then softly assigned to distinct Gaussian components via a probabilistic one-hot representation. This dual encoding strategy effectively captures both normalized numerical information and probabilistic cluster associations, improving the stability and efficacy of feature representation during generative model training. In the dis-dataset, we replace the BGMM-based encoding with fixed binning. All continuous features are discretized into categorical bins (e.g., cell lengths into 100 bins between 0–35 Å), and the resulting values are one-hot encoded. This yields a fully categorical representation for the entire dataset.
After transformation, the table expands to 525 columns in the cont-dataset and 2332 columns in the dis-dataset. These two encoding schemes allow us to benchmark model performance under both mixed-type and fully discrete data conditions. Next, we explored two generative models on both dis- and cont-type datasets to learn the underlying distribution of crystal symmetry data. The models are trained to synthesize new rows of tabular features that correspond to novel crystal structures that satisfy crystallographic constraints.
@@
For more details regarding GAN and VAE models, please refer to the Supplementary Materials Section I and Figs. S1 and S2). In a typical experiment, we train each GAN or VAE model with 250-1000 epochs, which is sufficient to converge the loss function values. Then we take the epoch that achieves the lowest loss function value to generate 100k - 500k samples for further consideration.
@@ III.3 Structure Relaxation to Targeted Local Environment
To represent the local atomic environment, the SO(3) descriptor $P$, with the option to include the explicit radial distribution function, is calculated to extract the radial and angular features of the neighbor density function Yanxon et al. (2020). For the purpose of this work, we selected a sp2 carbon (i.e., one centered atom is connected to 3 neighboring carbon with 1.42 Å  bond length and 120 bond angles in the same plane) to compute its power spectrum $P_{\text{ref.}}$ up to a radial cutoff of 2.1 Å. For each generated structure, we first evaluated the corresponding power spectrum $P$ for each atom, and then sought to minimize the summed mean squared error (MSE) loss function as follows:
@@
During optimization, symmetry constraints were naturally enforced (due to the use of reduced variables) to ensure chemical realism and structural validity. Combining with the subgroup representation, this descriptor-guided strategy not only shortens the computational overhead associated with iterative energy evaluations but also allows the system to escape undesirable local minima associated with conventional force-based optimization. While we may use the subgroup symmetry as the initial guess, it is important to note that the low-symmetry structure after relaxation does not necessarily evolve to high symmetry. This is because many pathways may exist to stabilize the Wyckoff site to a target reference environment. For more details, please refer to Fig. S3 and a motivating example in the Supplementary Section III.